OpenStudy (anonymous):

product of 3 numbers in G.P = 216 and the sum of their squares = 189 Then one of the numbers of this set can be?

OpenStudy (anonymous):

Let the three numbers be a/r , a , ar

OpenStudy (anonymous):

a/r * a * ar = a^3 = 216 a = 6

OpenStudy (shubhamsrg):

thats alright,,what eqn you made next?

OpenStudy (anonymous):

a^2/r^2 + a^2 + a^2 r^2 = 189

OpenStudy (shubhamsrg):

so?? what next/?

OpenStudy (experimentx):

put the value of a ... the equation is Quadratic in r^2

OpenStudy (anonymous):

36 r^4 + 36 = 153r^2

OpenStudy (experimentx):

solve for r^2

OpenStudy (anonymous):

36(r^4 + 1) = 153 r^2 r^4 + 1 = 4.25 r^2

OpenStudy (anonymous):

let r^2 = y y^2 + 1 = 4.25 y

OpenStudy (anonymous):

y^2 - 4.25y + 1 =0 y = 4 y = 0.5

OpenStudy (anonymous):

@experimentX any other easy way to solve this

OpenStudy (experimentx):

this is the easiest way!!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r = 2

OpenStudy (anonymous):

6 * 2 = 12

OpenStudy (anonymous):

very smart, easier i think than the idea of starting with \(a, ar, ar^2\) and then \(a^3r^2=216\)

OpenStudy (shubhamsrg):

well if you let 3 nos. to be a,b,c, we have abc = 216 ac=b^2 a^2 + b^2 +c^2 =189 from 1st and 2nd, b=6 and from 3rd, a^2 + c^2 + ac = 189 we know ac=36 from 1st eqn since b=6 (a+c)^2 - ac = 189 and (a-c)^2 + 3ac = 189 from these 2 eqns, a+c = 15 and a-c = 9 a= 12 and c=3 so nos are 12,6 and 3

OpenStudy (anonymous):

this is lot more simple thxxx